OpenStudy (anonymous):

Find the solutions of the inequality. | x - 5 | ≥ 3

myininaya (myininaya):

\[|f(x)| \ge 3 => f(x) \ge 3 \text{ or } f(x) \le -3 \] Think about it : Say we call this line f(x) -----|-----------|-------- -3 3 We want the distance to be greater than 3 That means we want f(x) to be ******] [**********

myininaya (myininaya):

Since f(x)=x-5 we have \[x-5 \ge 3 \text{ or } x-5 \le -3 \] You need to solve both of these inequalities for x

OpenStudy (anonymous):

first you set up both equation x-5≥ 3 x-5≥ -3

OpenStudy (anonymous):

then you bring over the 5 to the 3 and -3 so +5 and +5

OpenStudy (anonymous):

 I believe this will help you understand